1. What is the maximum value for the function y= 1/3sin5x
a. 1/3
b. 1/5
c. -1/3
d. -5
2. if g(x)=4sin 5x+k, what is the minimum value of g(x)?
a. 4 - k
b. k - 5
c. k
d. k - 4

Answer by alicealc(293)

The maximum value for sin ax is 1
the minimum value for sin ax is -1

so the answers:
1. the maximum value for y = 1/3 sin 5x is
1/3 * 1 = 1/3 (a)

2. the minimum value for g(x) = 4sin 5x + k is
4*(-1) + k
= -4 + k
= k - 4 (d)